A porcelain vase weighing 49.0N is supported by a tripod. The legs of the tripod are spaced at 120° apart and are inclined 60° with the horizontal. Find the force exerted by each leg of the tripod.
1
Expert's answer
2020-10-29T07:02:59-0400

The forse that acts into the first leg (120"^0" with the horizontal).

"F_1=Pcos 120^0=-49*0.5=-24.5" N


into the second leg (60"^0" with the horizontal):

"F_2=Pcos60^0=49*0.5=24.5" N


So. these forces are equal, but act into the opposite sides. The reason of it is the angle.